# The FinBen: An Holistic Financial Benchmark for Large Language Models

### v1

@@ D.2 Financial Evaluation Benchmarks
Financial evaluation benchmarks, such as the pioneering FLUE Shah et al. (2022), have been introduced to measure model performance in the financial sector, covering five key NLP tasks: financial sentiment analysis Shah et al. (2022), news headline classification Sinha and Khandait (2020), named entity recognition (NER) Salinas Alvarado et al. (2015), structure boundary detection and question answering (QA) Maia et al. (2018b). Building upon FLUE, FLARE (Xie et al., 2023b) added the evaluation of time-series processing capabilities, i.e., forecasting stock price movements. In addition, in Chinese financial benchmarks, there are more recently released chinese datasets like CFBenchmark Lei et al. (2023), DISC-FINSFT Chen et al. (2023b), and CGCE Zhang et al. (2023b). However, these benchmarks have a limited scope and have not yet to address more complex financial NLP tasks such as event detection Zhou et al. (2021), and realistic financial tasks, despite the fact that there were previous efforts on stock trading Liu et al. (2022); Han et al. (2023a, b).
@@ Appendix E Trading Accumulative Returns
Table 8 and below Figures show detail trading performance,
@@
Ticker Model CR (%) SR DV (%) AV (%) MD (%)
TSLA Buy and Hold -25.2137 -0.7203 4.4099 70.0043 57.6765
GPT-4 68.3089 2.8899 2.9780 47.2739 10.7996
GPT3.5-Turbo 25.2137 0.7203 4.4099 70.0043 51.3186
llama2-70B -31.4144 -1.0412 3.8014 60.3450 48.6173
gemini -0.3790 -0.0148 3.2271 51.2280 35.6707
NFLX Buy and Hold 34.6251 1.3696 3.1852 50.5634 20.9263
GPT-4 36.4485 2.0088 2.2860 36.2894 15.8495
GPT3.5-Turbo 7.9337 0.4610 2.1680 34.4160 17.9578
llama2-70B 33.8460 1.4741 2.8928 45.9216 20.3910
gemini 11.6298 1.0073 1.4546 23.0906 16.5106
AMZN Buy and Hold -16.4428 -0.7448 2.7812 44.1508 33.8847
GPT-4 10.5539 0.4923 2.7012 42.8802 22.9294
GPT3.5-Turbo 19.9636 0.9611 2.6171 41.5454 19.2191
llama2-70B 8.3595 1.9715 0.5342 8.4804 0.0000
gemini -2.3838 -0.5321 0.5645 8.9605 6.4291
MSFT Buy and Hold 17.2161 0.9709 2.2339 35.4623 15.0097
GPT-4 25.7826 1.5818 2.0535 32.5989 14.9889
GPT3.5-Turbo 20.4179 1.3600 1.8915 30.0259 20.3211
llama2-70B 27.7664 1.5708 2.2270 35.3524 15.0097
gemini 21.5082 1.3701 1.9777 31.3957 17.5051
COIN Buy and Hold -18.4787 -0.3369 6.9098 109.6904 60.5084
GPT-4 25.7631 0.5619 5.7761 91.6934 35.7526
GPT3.5-Turbo 25.1141 0.4772 6.6312 105.2669 53.9628
llama2-70B 15.1836 0.4395 4.3528 69.0979 35.3249
gemini 89.4782 1.7648 6.3879 101.4048 40.3246
AAPL Buy and Hold 12.7371 0.7759 2.0682 32.8323 20.6591
GPT-4 21.2334 1.9274 1.3879 22.0328 6.4237
GPT3.5-Turbo -27.0152 -1.9193 1.7734 28.1517 33.1619
llama2-70B 11.4855 1.1550 1.2529 19.8885 9.2776
gemini -5.3097 -0.3637 1.8392 29.1971 26.6450
GOOG Buy and Hold 6.3107 0.3081 2.5806 40.9660 21.1907
GPT-4 13.2811 0.9667 1.7308 27.4762 12.2209
GPT3.5-Turbo 0.9990 0.0614 2.0490 32.5265 20.9316
llama2-70B 17.0030 1.1057 1.9374 30.7546 13.2088
gemini 38.7956 3.0341 1.6110 25.5732 13.7311
NIO Buy and Hold -49.4263 -1.1895 5.2351 83.1048 52.2083
GPT-4 24.7684 0.9438 3.3063 52.4861 29.3384
GPT3.5-Turbo -28.9321 -1.0096 3.6105 57.3149 39.5907
llama2-70B -49.6947 -2.7868 2.2466 35.6639 42.6221
gemini 14.5673 0.6212 2.9543 46.8977 23.0110

### v2

@@ D.2 Financial Evaluation Benchmarks
Financial evaluation benchmarks, such as the pioneering FLUE (Shah et al., 2022), have been introduced to measure model performance in the financial sector, covering five key NLP tasks: financial sentiment analysis (Shah et al., 2022), news headline classification (Sinha and Khandait, 2020), named entity recognition (NER) (Salinas Alvarado et al., 2015), structure boundary detection and question answering (QA) (Chen et al., 2022a). Building upon FLUE, FLARE (Xie et al., 2023b) added the evaluation of time-series processing capabilities, i.e., forecasting stock price movements. In addition, in Chinese financial benchmarks, there are more recently released Chinese datasets like CFBenchmark (Lei et al., 2023), DISC-FINSFT (Chen et al., 2023b), and CGCE (Zhang et al., 2023b). However, these benchmarks have a limited scope and have not yet addressed more complex financial NLP tasks such as event detection (Zhou et al., 2021), and realistic financial tasks, despite the fact that there were previous efforts on stock trading (Liu et al., 2022; Han et al., 2023a, b).
@@ Appendix E Trading Accumulative Returns
Table 9 and the below Figures show detailed trading performance,
Ticker Model CR (%) SR DV (%) AV (%) MD (%)
TSLA Buy & Hold -25.2137 -0.7203 4.4099 70.0043 57.6765
GPT-4 68.3089 2.8899 2.9780 47.2739 10.7996
GPT-4o -0.8789 -0.0321 3.4531 54.8156 44.6842
GPT3.5-Turbo 25.2137 0.7203 4.4099 70.0043 51.3186
llama2-70B -31.4144 -1.0412 3.8014 60.3450 48.6173
llama3-70B -16.4424 -0.4847 4.2743 67.8519 55.5486
gemini -0.3790 -0.0148 3.2271 51.2280 35.6707
NFLX Buy & Hold 34.6251 1.3696 3.1852 50.5634 20.9263
GPT-4 36.4485 2.0088 2.2860 36.2894 15.8495
GPT-4o 5.5829 0.2592 2.7132 43.0702 17.4715
GPT3.5-Turbo 7.9337 0.4610 2.1680 34.4160 17.9578
llama2-70B 33.8460 1.4741 2.8928 45.9216 20.3910
llama3-70B 21.7374 0.9513 2.8788 45.6989 21.3478
gemini 11.6298 1.0073 1.4546 23.0906 16.5106
AMZN Buy & Hold -16.4428 -0.7448 2.7812 44.1508 33.8847
GPT-4 10.5539 0.4923 2.7012 42.8802 22.9294
GPT-4o 11.3626 0.7334 1.9520 30.9864 19.5964
GPT3.5-Turbo 19.9636 0.9611 2.6171 41.5454 19.2191
llama2-70B 8.3595 1.9715 0.5342 8.4804 0.0000
llama3-70B 11.1479 0.5405 2.5986 41.2509 28.2174
gemini -2.3838 -0.5321 0.5645 8.9605 6.4291
MSFT Buy & Hold 17.2161 0.9710 2.2339 35.4623 15.0097
GPT-4 25.7826 1.5818 2.0535 32.5989 14.9889
GPT-4o -5.3731 -0.5209 1.2997 20.6317 18.8223
GPT3.5-Turbo 20.4179 1.3600 1.8915 30.0259 20.3212
llama2-70B 27.7664 1.5708 2.2270 35.3524 15.0097
llama3-70B 21.1983 1.2628 2.1149 33.5724 15.0097
gemini 21.5081 1.3701 1.9777 31.3957 17.5051
AAPL Buy & Hold 12.7371 0.7759 2.0682 32.8323 20.6590
GPT-4 21.2335 1.9274 1.3879 22.0328 6.4237
GPT-4o -6.7540 -0.5693 1.4948 23.7285 20.7600
GPT3.5-Turbo 0.7110 0.0758 1.1817 18.7581 6.0818
llama2-70B 11.4856 1.1550 1.2529 19.8885 9.2776
llama3-70B -16.0835 -1.1985 1.6907 26.8394 25.9520
gemini 18.1718 1.7214 1.3300 21.1134 9.6467
GOOG Buy & Hold 6.3107 0.3081 2.5806 40.9660 21.1907
GPT-4 13.2811 0.9667 1.7308 27.4762 12.2209
GPT-4o 16.5072 1.0654 1.9520 30.9872 11.8863
GPT3.5-Turbo 0.9990 0.0614 2.0490 32.5265 20.9316
llama2-70B 17.0030 1.1057 1.9374 30.7546 13.2088
llama3-70B 17.5630 0.8872 2.4942 39.5934 19.2783
gemini 38.7956 3.0341 1.6110 25.5732 13.7311
DIS Buy & Hold -$0$0.0037 2.3667 37.5695 22.7722
GPT-4 31.3383 2.3931 1.6498 26.1904 12.3417
GPT-4o -20.2500 -1.3737 1.8573 29.4830 27.0246
GPT3.5-Turbo -7.1533 -0.5109 1.7641 28.0048 20.4278
llama2-70B -3.8257 -1.4323 0.3365 5.3420 4.1451
llama3-70B -25.5829 -1.5579 2.0690 32.8437 31.3391
gemini 8.6692 0.8015 1.3627 21.6321 18.4815
GM Buy & Hold 0.3393 0.0179 2.3823 37.8181 23.0317
GPT-4 10.5648 0.7671 1.7351 27.5443 11.1285
GPT-4o -7.0147 -0.5263 1.6792 26.6569 21.5978
GPT3.5-Turbo -17.6385 -0.9692 2.2928 36.3976 23.0317
llama2-70B 8.4911 2.6369 0.4057 6.4402 2.1318
llama3-70B 25.9335 1.9823 1.6483 26.1657 13.2485
gemini 18.6257 2.4672 0.9511 15.0989 3.0369
NIO Buy & Hold -49.4263 -1.1895 5.2351 83.1048 52.2083
GPT-4 24.7684 0.9438 3.3063 52.4861 29.3384
GPT-4o -48.3748 -1.5026 4.0562 64.3897 59.4037
GPT3.5-Turbo -28.9321 -1.0096 3.6105 57.3149 39.5907
llama2-70B -49.6947 -2.7868 2.2466 35.6639 42.6221
llama3-70B -28.6668 -0.7094 5.0912 80.8202 37.1544
gemini 14.5673 0.6212 2.9543 46.8977 23.0110
COIN Buy & Hold -18.4787 -0.3369 6.9098 109.6904 60.5084
GPT-4 25.7631 0.5619 5.7761 91.6934 35.7526
GPT-4o -14.2451 -0.2892 6.2049 98.4997 65.3090
GPT3.5-Turbo 25.1141 0.4772 6.6312 105.2669 53.9628
llama2-70B 15.1836 0.4395 4.3528 69.0979 35.3249
llama3-70B 19.8876 0.3749 6.6842 106.1076 55.7225
gemini 89.4782 1.7648 6.3879 101.4048 40.3246
@@
Figure 10: Accumulative Returns of LLM Trading Strategies on DIS
Figure 11: Accumulative Returns of LLM Trading Strategies on GM
Appendix F FinLLM challenge
Based on our proposed FinBen, we organized the FinLLM Share Task during the FinNLP-AgentScen Workshop at IJCAI 2024999https://sites.google.com/nlg.csie.ntu.edu.tw/finnlp-agentscen/shared-task-finllm?authuser=0, known as the FinLLM Challenge. This challenge not only tests the abilities of LLMs but also promotes ongoing research into their application within the financial sector, highlighting FinBen’s critical contribution to the advancement of financial analytics.
The FinLLM Challenge is a specialized shared task tailored for LLMs, targeting a comprehensive range of financial problems through three subtasks: financial classification, financial text summarization, and single stock trading. To rigorously evaluate the capabilities of financial LLMs, we have curated three distinct datasets corresponding to each of these subtasks, as detailed in Table 10. This structured approach ensures a holistic and effective assessment of LLM performance across diverse financial scenarios.
Table 10: Tasks and Datasets of FinLLM Challenge.
Category Tasks Datasets Evaluation Metrics
Training set Test set
Task 1 Financial Classification 7.75k 969 F1 Score, Acc
Task 2 Financial Text Summarization 8k 2k ROUGE-1, ROUGE-2, ROUGE-L, BERTScore
Task 3 Single Stock Trading 291 225 Sharpe Ratio, Cumulative Return,
Maximum Drawdown, Daily and Annualized Volatility,
F.1 Tasks and Datasets
Task 1: Financial Classification. This task, inherited from FinBen’s financial classification task, focuses on argument unit classification to test the capabilities of LLMs to identify and categorize texts as premises or claims. It consists of 7.75k training data and 969 test data to categorize sentences as claims or premises. We use two metrics to evaluate classification capability, like F1 and Accuracy. F1 score is used as the final ranking metric.
Task 2: Financial Text Summarization. This task, inherited from FinBen’s generation task, is designed to test the capabilities of LLMs to generate coherent summaries. It provides 8k training data and 2k test data for abstracting financial news articles into concise summaries. We utilize three metrics, such as ROUGE (1, 2, and L) and BERTScore, to evaluate generated summaries in terms of Relevance. ROUGE -1 score is used as the final ranking metric.
Task 3: Single Stock Trading. This task, inherited from FinBen’s Trading task, aims to evaluate LLMs’ ability to make sophisticated decisions in trading activities, which is currently restricted by human’s limited ability to process large volumes of data rapidly. It specifically provides 291 data different from FinBen datasets, to evaluate LLMs on sophisticated stock Decisions. We offer a comprehensive assessment of profitability, risk management, and decision-making prowess by a series of metrics, such as Sharpe Ratio (SR), Cumulative Return (CR), Daily (DV) and Annualized volatility (AV), and Maximum Drawdown (MD). Sharpe Ratio (SR) score is used as the final ranking metric.
F.2 Model Cheating Detection
To measure the risk of data leakage from the test set used in training, we introduce the Data Leakage Test (DLT). The DLT calculates the difference in perplexity between the training set and the test set. A larger difference indicates a lower likelihood of model cheating, while a smaller difference suggests a higher likelihood. For our FinLLM Challenge, we invite Top-3 participant teams per task for cheating detection.
F.3 Participants and Automatic Evaluation
There are 35 teams registered for FinLLM Challenge, with 11 teams submitting their system description papers. Participants can opt to join one or more task(s).
As shown in Table 11, the top 3 teams achieved outstanding performance in Task 1. Their models’ F1 scores were comparable to LlaMA3-8B, although slightly inferior to GPT-4 and LLaMA2-70B, yet significantly outperformed FinMA and other models. The results in Table 11 further demonstrate that our FinLLM share task provides an excellent framework for participating teams to achieve superior experimental outcomes.
Table 11: The Result of Taks 1: Financial Classification
Teams ACC F1 MCC
Team Barclays 0.7626 0.5237 0.7427
Albatross 0.7574 0.5174 0.7555
L3iTC 0.7544 0.5149 0.7581
Wealth Guide 0.7513 0.5018 0.7406
Finance Wizard 0.7286 0.4554 0.7008
CatMemo 0.711 0.4199 0.6818
Upaya 0.709 0.4166 0.6941
Vidra 0.7079 0.4141 0.69
jt 0.4933 0.0141 0.5905
As illustrated in Table 12, in terms of the Rouge-1 metric, the models of these three teams surpassed all other models, demonstrating superior performance. The results in Table 2 indicate that, for financial generation tasks, our provided dataset and model framework help participating teams leverage their strengths and achieve better outcomes.
Table 12: The Result of Taks 2: Financial Text Summarization
Teams Rouge-1 Rouge-2 Rouge-L BertScore BartScore
Wealth Guide 0.308893532 0.179468097 0.281924302 0.85959909 -4.961457408
Albatross 0.369077581 0.201058395 0.322684316 0.872049115 -3.933526929
LBZ 0.534616211 0.358105428 0.492179554 0.911732047 -3.407560172
L3iTC 0.366093426 0.187210467 0.304610677 0.875037043 -4.257126737
Finance Wizard 0.521037018 0.34060938 0.473530112 0.90836845 -3.497988865
Vidra 0.284955468 0.134760859 0.228638961 0.858682767 -4.169740305
Revelata 0.500411369 0.333023818 0.464356474 0.907018743 -3.805486962
Upaya 0.529459817 0.358203218 0.486046685 0.910644962 -3.45155009
As shown in Table13, the Top-1 Wealth Guide team excelled in the Sharpe Ratio metric, surpassing other teams and demonstrating outstanding performance. While it may not match the performance of GPT-4, it still outperforms other large models. These results from Table 3 once again underscore the significance of organizing the FinLLM share task. The FinLLM Challenge not only assesses the performance of large language models (LLMs) but also fosters further research into applying LLMs in the financial domain.
Table 13: The Result of Taks 3: Single Stock Trading
Teams Sharpe Ratio Sharpe Ratio-DRIV Sharpe Ratio-FORM Sharpe Ratio-JNJ Sharpe Ratio-MSFT
Wealth Guide 0.9263852228 0.485625528 1.585611423 0.078737051 1.555566991
Upaya 0.467489019 0.380232272 0.108506918 -1.102831656 -0.278385232
Albatross 0.48383204 0.251306057 -1.435471054 -1.558522674 1.309971626
CatMemo -0.619939784 -1.393291177 0.175932289 0.383243051 -0.879157198
Appendix G Performances of non-LLM methods
In this section, we present the performances of non-LLM methods on stock movement prediction and financial NLP tasks from previous papers. Note that non-LLM methods are task-oriented, each model can only run on a specific task.
G.1 Stock movement prediction
Stock movement prediction performance of non-LLM models are shown in Table 14. The results are from (Xie et al., 2023b).
Table 14: Stock movement prediction performance of non-LLM models, measured with the accuracy (ACC) and the Matthews correlation coefficient (MCC). The best performance is in bold.
Method BIGDATA22 ACL18 CIKM18
ACC MCC ACC MCC ACC MCC
Logistic regression (LR) 0.53 0.02 0.52 0.04 0.53 -0.04
Random forest (RF) 0.47 -0.11 0.52 0.03 0.54 0.01
LSTM 0.51 0.01 0.53 0.06 0.53 0.02
Attention LSTM (ALSTM) 0.49 -0.03 0.52 0.04 0.53 -0.01
Adv-ALSTM 0.50 0.01 0.53 0.07 0.54 0.02
DTML 0.52 0.07 0.58 0.18 0.54 -0.00
XGBoost 0.52 -0.04 0.49 -0.02 0.58 0.07
XGBRegressor 0.46 -0.13 0.50 -0.01 0.53 -0.03
ALSTM-W 0.48 -0.01 0.53 0.08 0.54 0.03
ALSTM-D 0.49 0.01 0.53 0.07 0.50 -0.04
StockNet 0.53 -0.00 0.54 -0.03 0.52 -0.02
SLOT 0.55 0.10 0.59 0.21 0.56 0.09
G.2 Financial NLP tasks
BERT-based model results of financial NLP tasks are shown in Table 15. The results are from (Shah et al., 2022).
Table 15: Financial NLP tasks performances of BERT-based models. The best performance is in bold.
Method FPB Headline NER FiQA SA
Accuracy AvgF1 F1 MSE
BERT-base 0.856 0.967 0.79 0.073
FinBERT 0.872 0.968 0.8 0.070
FLANG-BERT 0.912 0.972 0.83 0.054
ELECTRA 0.881 0.966 0.78 0.066
FLANG-ELECTRA 0.919 0.98 0.82 0.034
Limitations
Despite the novel efforts to benchmark LLMs in the financial domain through FinBen, we acknowledge several inherent limitations that could impact the benchmark’s effectiveness and applicability:
Despite the novel efforts to benchmark LLMs in the financial domain through FinBen, several limitations could impact its effectiveness and applicability: Dataset Size Limitations: The restricted size of available datasets, a common issue in open-source financial data, may affect the models’ financial understanding and generalization across various contexts. Model Size Limitations: Due to computational constraints, our evaluation was limited to the LLaMA 70B model, potentially overlooking the capabilities of larger or differently architected models. Generalizability: The tasks, particularly trading and forecasting, are based on American market data and English texts, possibly limiting the benchmark’s applicability to global financial markets. Potential Negative Impacts: While FinBen aims to advance financial language understanding, it is crucial to consider potential misuse, such as propagating financial misinformation or exerting unethical influence on markets. Responsible usage and further safeguards are essential [^10] For a detailed ethical and legal statement concerning this work, please see Appendix..
Ethical Statement
The development and dissemination of the FinBen by the authors carry full responsibility for any potential violation of rights or arising legal issues. All raw data we used are publicly available and do not contain any personal information. Diligent efforts have been undertaken to ensure the construction of the FinBen respects privacy and conforms to established ethical guidelines. The datasets compiled within FinBen are shared under the MIT license, with the expectation that users agree to adhere to its conditions.